Status Report

NASA Space Station On-Orbit Status 1 Feb 2004

By SpaceRef Editor
February 2, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 1 Feb 2004

All ISS systems continue to function nominally, except as noted previously or below.    MCC-Houston today observed a moment of silence at 9:16am EST in remembrance of the Columbia crew, who died a year ago.  Opening ceremonies for the Superbowl also included a special tribute to the crew.  “The cause in which they died will continue.  Our journey into space will go on.”  (President George W. Bush, 2/1/03)

After wake-up (1:00am EST) and breakfast, CDR Michael Foale activated the MSG (Microgravity Science Glovebox) for today’s PromISS (study of fundamental processes underlying protein crystallization) activities.

Yesterday’s Progress 13P (M1-11 or No. 260) opening activities went without issues.  The crew performed atmospheric sampling via IPD Draeger tubes before accessing the vestibule between the Service Module (SM) and 13P.  No ammonia from the HEAT hardware was detected by the crew via the IPD sampler or by smell.  After waiting for equalization of pressures between the station and the Progress, the crew opened the Progress hatch and was greeted with the welcome smells of fresh fruit.  The ISS total atmospheric pressure was raised by approximately 3 mmHg as a result of the equalization with the Progress.

Today, the crew the crew removed the SSVP probe-and-cone docking mechanism to clear the passage for the ensuing transfer operations.   [The SSVP is the “classic” probe-and-cone type, consisting of an active docking assembly (ASA) with a probe (SSh), which fits into the cone (SK) on the passive docking assembly (PSA).  The ASA is mounted on the Progress’ cargo module (GrO), while the PSA sits on the docking ports of the SM, FGB and DC-1.]

Afterwards, Mike Foale and Alexander Kaleri began with Progress 13P unloading and cargo transfers, to continue over the next few days.  An updated list was uplinked a few days ago to support the transfer and stowage operations with the help of the IMS (inventory management system).

Wearing protective gear (face masks, protective goggles, surgical gloves) the crew conducted the first hardware monitoring and photo imagery of the JAXA (Japanese Space Exploration Agency)-sponsored GCF (Granada Crystallization Facility), which they first transferred to the CGBA (Commercial Generic Bioprocessing Apparatus).   [The experiment, which is classified as toxicity level 1, deals with the growth of macromolecular crystals from solution with higher performance and reliability in terms of crystal quality and/or crystal size, compared to those grown in earth-based biochemistry laboratories and crystal growth laboratories.  It needs to be maintained at 20 deg C and should never exceed 25 deg C.  After a visual inspection of the three GCF-02 crystallization units inside the powered-down CGBA, the crew took shots of them with the Nikon D1 digital camera for subsequent downlink to MCC-M/TsUP via OCA, preparatory to subsequent power-up of the CGBA by Mike Foale, to keep it at 20 degC.  The GCF-01 transport container was discarded.]

Working at the MSG, Science Officer Foale continued PromISS-3 hardware installation and activation, which he had started on 1/30 with set-up part 1.   [PromISS-3 activation activity was marked time critical so that the start of the crystallization process was captured via video.]

Sasha Kaleri transferred the new samples of the Russian technology experiments BTKh-2 (Mimetic-K) and BTKh-20 (Interleikin) from Progress 13P to the SM and activated the hardware.  [Both experiments deal with growing protein crystals with high-quality crystalline structure.  The samples were transferred from Progress in the Luch-2 Kit (a biocrystallizer assembly comprising four multi-purpose Luch-2 biological crystallization cartridges {UBK}), for installation in the on-board AQUA-01 container during crystallization.]

Afterwards, the FE continued integration of the Progress 13P into the ISS/SM systems, today installing the LKT local temperature sensor commutator (TA251MB) of the BITS2-12 onboard telemetry system and its ROM (read-only memory) unit (PZU/TA765B) in the cargo module, a one-hour task.

Kaleri conducted the daily routine maintenance of the SM’s SOZh life support system and collected the weekly data readings of the SM toilet flush counter readings (with inspection of the SP urine collection and pretreat assembly) and SVO water supply status counter readings, both for calldown to TsUP.

The crew worked out on TVIS treadmill, RED exerciser and VELO cycle with load trainer.

A summary was uploaded of the upcoming new TVIS loads test on 2/11, which will use IWIS (internal wireless instrumentation system), Russian sensors over Russian ground sites, and fixed video cameras.  [The test is to get structural loads data on the crew’s maximum TVIS harness loading, before the long-planned TVIS gyro repair is attempted in March, just in case the gyro repair is unsuccessful.]

The station is currently again in solar-oriented XPOP (x-axis perpendicular to orbit plane) attitude, until 2/5.

Today’s CEO (crew earth observations) targets, in the current XPOP attitude constrained by flight rule to fewer near-vertical targets due to Lab window shutter closure and current condensation-prevention plan (limited to 90 min. in 24 hours), were Cyclone Frank, Indian Ocean (looking right of track for this strong category 3 storm.  The cloud mass is 1-2 deg off track, the eye ~6.5 degrees off track), Buenos Aires, Argentina (looking nadir and right of track), and Tuamotu Islands, Pacific (400mm-lens:  ISS passed over the eastern sector of this twin island chain.  Cloud approached from the W: looking right (E) for coral-fringed atolls for the global coral-mapping project).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

SpaceRef staff editor.